# Emergence Learning: A Rising Direction from Emergent Abilities and a Monosemanticity-Based Study

### 17 Dec 2023 v1

@@
Abstract
In the past 20 years, artificial neural networks have become dominant in various areas, continually growing in scale. However, the current analysis of large models has mainly focused on functionality, overlooking the influence of scale differences on their properties. To address this, we propose the concept of Emergence Learning, which emphasizes the significance of scale. By studying models of different scales, we have identified a key factor in achieving higher performance in large models: the decrease of monosemantic neurons. Building on this insight, we propose a proactive approach to inhibit monosemanticity for improved performance. Our solution involves a two-phase process that includes monosemantic neuron detection and inhibition, supported by theoretical analysis. Experimental results on various tasks and neural networks demonstrate the effectiveness of our proposed method.
Following the idea of Emergence Learning, though drawing inspiration from scaling phenomena, the applicability of our method is not restricted to large scale alone. Therefore, the experiment is self-contained. However, extending this research to very large-scale datasets is appealing yet impossible for research departments due to limited resources. We are delighted to share the first co-authorship and eagerly await collaboration from any AI company before submission.
Deep Learning, Emergence, Monosemantic
1 Introduction
Inspired by biological neurons, the artificial neural network (ANN) was proposed in the last century (S & Walter, 1943). However, its activity diminished for decades before experiencing great success after the 2010s (Krogh, 2008; Krizhevsky et al., 2012). One of the major differences compared to previous models is the relatively large scale.
@@ 1 Introduction
In recent years, large models have achieved remarkable results in various fields (Floridi & Chiriatti, 2020; Touvron et al., 2023). Compared to previous works, these models have much larger scales in terms of datasets, model sizes, and training quantities. Though the idea of “the larger, the better” is common sense, unexpected improvements could occur solely due to larger scales. Emergence, as an intriguing phenomenon in large-scale models, refers to the gradual improvement of model performance before the scale reaches a certain threshold, followed by a rapid enhancement once the threshold is surpassed (Wei et al., 2022). Increasing evidence suggests that the surprises we experience may not come from the functionality and mechanism, but rather from the underlying nature of scale changes. A critical question arises: people increase model scale, but increase for what?
Therefore, we propose Emergence Learning (EmeL for short), which emphasizes the property differences brought about by scale differences, and seeks to uncover potential principles to improve models.
Related to EmeL, recent research has also analyzed large models and observed changes resulting from differences in scale (Geva et al., 2021b; Gurnee et al., 2023). However, the research paradigm following smaller-scale models hinders a more in-depth investigation. A potential case we found is the study on monosemantic and polysemantic neurons. Through statistical analysis on the relationships between neurons and input features, a neuron is considered monosemantic if it forms a 1-1 correlation with its related input feature (Bau et al., 2020; Elhage et al., 2022) (Figure. 1(b)). In contrast, polysemantic neurons activate for several features that are weakly correlated (Trenton et al., 2023; Chris et al., 2020; Gabriel et al., 2021; Elhage et al., 2022) (Figure. 1(c)).
Figure 1: Demonstration of important concepts and motivation: (a) A multilayer neuron network for the image classification task. The middle layer is highlighted. (b) A monosemantic neuron (orange) ideally activates for one specific type of feature. (c) A polysemantic neuron activates for multiple features. (d) A distributed feature is classified based on the activation of multiple neurons. (e) ANNs could be similar to biological brains. Small-scale networks (both biological and artificial) cannot support complex functionality. (f) To solve a difficult problem that requires complex reasoning, pupils rely on rote memorization, while ANNs store QA pairs using monosemantic neurons as key-value pairs. (g) Large-scale networks can learn and master solving skills related to integration. (h) Adults and large-scale ANNs are capable of decomposing features and inferring answers using complex neuron circuits. This ability reduces reliance on rote memory and monosemanticity.
Recent studies have shown that monosemantic neurons are sparser in larger models and require more complex methods to detect (Gurnee et al., 2023; Trenton et al., 2023). Experimental results discussed in Appendix. A reveal more deficits in monosemantic neurons, e.g., when an input’s related monosemantic neuron is disabled, smaller models incur much larger errors compared to larger models (Gurnee et al., 2023). Those results indicate a positive relationship between the decrease of monosemanticity and the increase of model scale, which is highly likely to result in good performance.
However, following the favor of monosemantic neurons from studies on small-scale models, existing works focus on enhancing or extracting monosemanticity. In their work, Elhage et al. (2022) introduce the softmax linear unit to create monosemantic models. Additionally, Gurnee et al. (2023) modify the sparse probe by using multiple heads to classify neurons with decreasing monosemanticity. To provide interpretability in the feed-forward layer, key-value pairs are constructed, aiming to map each key to identifiable patterns (Geva et al., 2021b). Following the paths of small-scale models, these works overlook the potential for improvement direction behind the harmfulness of monosemanticity.
Therefore, we propose our hypothesis and question:
  • Hypothesis: As the model scale increases, the decrease of monosemantic neurons is a key factor in achieving higher performance.
  • Question: Can we proactively induce the decrease of monosemantic neurons to achieve high performance?
An example is given below to illustrate our idea from the angle of EmeL.
  • From the perspective of scale differences, when the model is small, the limited number of neurons makes it difficult to abstract and analyze the input, and to provide functional support for integrating a large amount of scattered information (Figure. 1(e)). To achieve good results, neurons tend to specialize in certain features, resulting in stronger monosemanticity (Figure. 1(f)).
    Similarly, we can analogize large models to adults, allowing them to learn and master solving skills related to integration (Figure. 1(g)). This is because adults have accumulated knowledge and have a more mature brain development, similar to the huge training volume and parameter size of large models. In this process, adults gradually reduce the reliance on rote memorization and improve their reasoning abilities for new problems (Figure. 1(h)). This can explain why large models have fewer monosemantic neurons and stronger generalization abilities in complex tasks.
To actively guide the reduction of monosemanticity, we propose a 2-phase solution: (i) monosemantic neuron detection and (ii) monosemantic neuron inhibition. Existing works build feature datasets and detect neuron activation when input contains specific features to find monosemantic neurons (Gurnee et al., 2023; Trenton et al., 2023; Steven et al., 2023b). However, these methods have efficiency and flexibility issues. Instead, we provide an effective online metric for detection. To inhibit monosemantic neurons, we identify the drawbacks of simple methods and propose a theoretically supported Reverse Deactivation method to modify monosemantic neurons, guiding the model to reduce monosemanticity during training. Our approach can be integrated as a parameter-free and flexible insertion module called MEmeL, which can adapt to various neural network structures and introduce no computational overhead during testing.
@@
To validate the effectiveness and generalizability of our method, we conducted tests on language, image, and physics simulation tasks, applying MEmeL to milestone models such as Bert and ConvRNN. The architectures covered Transformer, CNN, and RNN. The experimental results demonstrated the effectiveness of our method.
We summarize our contribution as follows:
  • We raise the concept of “Emergence Learning”, though far behind solving it, referring to a more recent research direction aside in middle- to large-scale models, in which patterns of neurons hardly appear in small models. The target is to identify factors related to the difference of scales and design specific solutions to induce the occurrence of emergence or improve performance.
  • To validate our hypothesis that the decrease of monosemanticity could be a key factor in high performance, we propose an efficient and effective method to detect monosemanticity.
  • We design solution RD to inhibit monosemanticity with theoretical guarantee, which overcomes an inherent drawback of naive solutions.
  • We conduct experiments in various tasks with different types of neural networks to validate the effectiveness of our method.
    (We are exploring our computational resources to conduct larger-scale experiments.)
The rest of the paper is organized as follows: Section 2 provides the formal problem definition. Section 3 introduces our method for decreasing monosemanticity. Experimental results are presented in Section 4. We review related works in Section 5. Finally, we conclude our paper in Section 6.
2 Problem Definition
In this section, we will first introduce the basic concepts. Then, we will provide formal definitions for two problems towards our goal: the detection and inhibition of monosemantic neurons.
2.1 Basic Concepts
We define a deep learning model as $\mathcal{F}$, with input and output represented as $X$ and $Z$, respectively. The relationship between the two is given by the equation:
$Z=\mathcal{F}(X).$
The goal of training model $\mathcal{F}$ is to update its parameters so that the output $Z$ becomes closer to a provided or unknown ground truth $Y$.
Deep learning models are composed of multiple layers, with each layer containing multiple neurons. We denote a specific layer of neurons as $H=\{h_{1},h_{2},\cdots\}$. In addition to existing methods for model optimization, we considering two new targets during training to improve performance:
  • Detection of monosemantic neurons of $H$.
  • Inhibition of monosemantic neurons of $H$.
2.2 Monosemanticity Detection
To detect monosemantic neurons, existing methods are inflexible as they require a predefined and manually labeled feature dataset (Gurnee et al., 2023; Trenton et al., 2023). Moreover, detection methods like probes necessitate training and offline inference for statistical analysis, which can be costly for online training. To address these limitations, there is a need for a robust metric $\phi$ that can accurately detect monosemantic neurons. This metric should fulfill the following criteria:
  • Generality. Does not rely on any specific dataset.
  • Efficiency. Fast online detection during training.
2.3 Monosemanticity Inhibition
To inhibit a detected monosemantic neuron $h\in H$, we need to design the optimization strategy $\omega$ for $\mathcal{F}$. Without loss of generality, when we consider a particular layer of neurons $H$, we can divide the model $F$ into a frontal model $f_{1}$ and a followed model $f_{2}$, so that:
$H=f_{1}(X),Z=f_{2}(H,X).$
The problem can be solved in two phases:
  • Optimize the frontal model $f_{1}$:
    $\omega(f_{1})\rightarrow{f_{1}^{*}},$
    so that the neuron $h$ is less activated for input $X$:
    $\phi(h^{*})<\phi(h),$
    where $h\in H=f_{1}(X)$ and $h^{*}\in H^{*}=f_{1}^{*}(X)$.
  • Optimize the followed model $f_{2}$:
    $\omega(f_{2})\rightarrow{f_{2}^{*}},$
@@ 2.3 Monosemanticity Inhibition
    so that the output is still robust when neuron $h$ is deactivated. Formally, replace $h\in H$ with a weakly activated value $h^{\prime}$ to obtain $H^{\prime}$, we expect:
    $\mathcal{L}(f_{2}^{*}(H^{\prime},X))<\mathcal{L}(f_{2}(H^{\prime},X)),$
    where $\mathcal{L}$ is the loss function, $\phi(h^{\prime})<\phi(h)$.
@@
3 Methods
3.1 Metric for Monosemanticity
As we want to inhibit monosemanticity during training, designing a metric that is general for different tasks and efficient to calculate is important. In this subsection, we first give the definition of our proposed metric. Then we analyze its advantages.
3.1.1 Metric Definition
Formally, we define our metric Monosemantic Scale (MS for short):

### 15 Feb 2024 v2

# Learning from Emergence: A Study on Proactively Inhibiting the Monosemantic Neurons of Artificial Neural Networks

@@
Abstract.
Recently, emergence has received widespread attention from the research community along with the success of large language models. Different from the literature, we hypothesize a key factor that highly promotes the performance during the increase of scale: the reduction of monosemantic neurons that can only form one-to-one correlations with specific features. Monosemantic neurons tend to be sparser and have negative impacts on the performance in large models. Inspired by this insight, we propose an intuitive idea to identify monosemantic neurons and inhibit them. However, achieving this goal is a non-trivial task as there is no unified quantitative evaluation metric and simply banning monosemantic neurons does not promote polysemanticity in neural networks. Therefore, we propose to learn from emergence and present a study on proactively inhibiting the monosemantic neurons in this paper. More specifically, we first propose a new metric to measure the monosemanticity of neurons with the guarantee of efficiency for online computation, then introduce a theoretically supported method to suppress monosemantic neurons and proactively promote the ratios of polysemantic neurons in training neural networks. We validate our conjecture that monosemanticity brings about performance change at different model scales on a variety of neural networks and benchmark datasets in different areas, including language, image, and physics simulation tasks. Further experiments validate our analysis and theory regarding the inhibition of monosemanticity.
Deep Learning, Artificial Neural Networks, Emergent Ability, Monosemanticity
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; August, 2024; Spainisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Figure 1. Demonstration of important concepts:(a) A multilayer neuron network for the image classification task. The middle layer is highlighted. (b) A monosemantic neuron (orange) ideally activates for one specific type of feature. (c) A polysemantic neuron activates for multiple features. (d) A distributed feature is classified based on the activation of multiple neurons.
@@ 1. Introduction
The activity of artificial neural networks diminished for decades before experiencing great success after the 2010s (Krogh, 2008; Krizhevsky et al., 2012). One major difference compared to previous models is the increasing scale. In recent years, large neural networks have much larger scales in terms of datasets, model sizes, and training quantities, which have achieved remarkable results in various fields (Floridi and Chiriatti, 2020; Touvron et al., 2023). Given increasing scales, the performance improvements can usually be estimated under the scaling law (Kaplan et al., 2020), which follows stable but slow increasing trend of power laws (OpenAI, 2023). However, improvements that significantly deviate from estimation occur when the scale increases. Emergence, an intriguing phenomenon in large-scale models, refers to the gradual improvement of model performance before the scale reaching a certain threshold, followed by a rapid enhancement once the threshold is surpassed (Wei et al., 2022). Increasing evidence suggests that the surprises may not arise from new module and architecture designs, but rather from the underlying nature of scale changes (Zhou et al., 2023). A critical question arises: people increase the model scale and get better results, but what has changed underlying the process?
From the perspectives of monosemantic and polysemantic neurons, some pioneer works try to interpret the performance of small and large-scale models, respectively. Through statistical analysis of the relationships between neurons and input features, a neuron is considered monosemantic if it forms a 1-1 correlation with its related input feature (Bau et al., 2020; Elhage et al., 2022) (Figure 1-b). In contrast, polysemantic neurons activate for several features that are weakly correlated (Trenton et al., 2023; Chris et al., 2020; Gabriel et al., 2021; Elhage et al., 2022) (Figure. 1-c). Researchers find that smaller models incur larger errors (Gurnee et al., 2023) when an input’s related monosemantic neuron is disabled. Following the favor of monosemantic neurons from studies on small-scale models, existing works focus on enhancing or extracting monosemanticity (Gurnee et al., 2023; Elhage et al., 2022; Trenton et al., 2023). However, some evidences support that monosemantic neurons are sparser in larger models (Gurnee et al., 2023; Trenton et al., 2023). Those observations imply a positive relationship between the decrease in monosemanticity and the increase in model scale. In other words, monosemanticity could be negatively related to large-scale and good performance. In this paper, inspired by those observations, we propose an important hypothesis that the decrease of monosemantic neurons is a key factor in achieving higher performance as the model scale increases. Figure. 4 in Appx. A shows an example analogy to demonstrate our idea.
Since the research community does not realize the above hypothesis, we rather conclude the current paradigm of training neural networks as a passive process in decreasing monosemantic neurons. This raises an interesting question: can we proactively induce the decrease of monosemantic neurons in artificial neural networks to achieve high performance?
In this paper, we propose to learn from emergence and present a study on proactively inhibiting the monosemantic neurons, which is achieved in two phases:(i) monosemantic neuron detection and (ii) monosemantic neuron inhibition. Unfortunately, it is challenging to design detection and inhibition methods. First, strictly defining monosemantic neurons is still under discussion in quantitative analysis, i.e., detection is impossible without a consensual definition. Besides, existing works (Gurnee et al., 2023; Trenton et al., 2023; Steven et al., 2023a) for detecting the activation degree of neurons introduce extra computation and may suffer from the efficiency issue. Second, we prove that simply prohibiting the activation of monosemantic neurons will intensify the monosemanticity of artificial neural networks. To solve these challenges, we first propose a new metric to measure the monosemanticity of neurons with the guarantee of efficiency for online computation. Then, we identify the drawbacks of simple methods and propose a theoretically supported Reverse Deactivation method to suppress monosemantic neurons and promote the ratios of polysemantic neurons in training neural networks. Our contributions are summarized as follows:
  • Inspired by emergence, we propose a novel idea to study the impact of monosemantic and polysemantic neurons on the performance of artificial neural networks. Different from the literature, we propose a hypothesis that monosemanticity could be negatively related to good performance on large-scale models.
  • There is no quantitative definition of monosemanticity with high computational efficiency. To effectively detect monosemantic neurons, we propose a new evaluation metric of monosemanticity with a theoretical guarantee on online computation.
  • To overcome the inherent drawback of the naive inhibition methods, we propose a theoretically supported reverse deactivation method to suppress monosemantic neurons. It can be integrated as a parameter-free and flexible insertion module, which can adapt to various neural network structures and introduce no computational overhead during testing.
@@
  • To validate the effectiveness and generalizability of our method, we conducted tests on various tasks including language, image, and physics simulation. The proposed method MEmeL is applied to milestone models such as Bert, ConvRNN, and different architectures like Transformer, CNN, and RNN. The experimental results demonstrate that MEmeL performs well on different neural networks and corresponding tasks.
2. Preliminary
Given a training set $D=\{(\mathbf{x},\mathbf{y})\}$ of labeled examples, artificial neural networks aim to train a neural network model $f(\mathbf{W};\mathbf{x})$ parameterized by $\mathbf{W}$ so that the predicted output of $\mathbf{x}$ is close to the ground truth $\mathbf{y}$. Under the supervised learning task, we take the cross entropy function $\mathcal{L}(\mathbf{W};D)$ as the example:
(1) $\min_{\mathbf{W}}\mathcal{L}(\mathbf{W};D)=\min_{\mathbf{W}}-\sum_{(\mathbf{x},\mathbf{y})\in D}\mathbf{y}\log f(\mathbf{W};\mathbf{x}),$
where $\mathbf{x}=[x_{1},\cdots,x_{d}]\in\mathbb{R}^{d}$ is a $d$-dimensional vector representing the $n$-th input data sample, such as a weather record series covering $d$ days or an image with height $\times$ width $=d$ pixels.
Without loss of generality, let $f(\mathbf{W};\mathbf{x})$ be a $L$-th layer fully connected neural networks (Haykin, 1994). We denote $h_{i}^{\ell}$, $\sigma_{i}^{\ell}(\cdot)$, $z_{i}^{\ell}$ as the input after linear transformation, activation function, and output of the $i$-th hidden unit at $\ell$-th layer, respectively. Then, we may formally define the computation within neurons as follows:
(2) $h_{j}^{\ell}=\sum_{i}w_{ij}^{\ell}z_{i}^{\ell-1},$
(3) $z_{i}^{\ell}=\sigma_{i}^{\ell}(h_{i}^{\ell}),$
where $z_{i}^{0}\in\mathbf{z}^{0}=\mathbf{x}$, $w_{ij}^{\ell}$ denotes the weight of $i$-th hidden neuron at $(\ell-1)$-th layer to $j$-th hidden neuron at $\ell$-th layer. We use $\mathbf{o}=\mathbf{z}^{L}$ to denote the model output. Existing works on monosemanticity mainly study the layers of values outputted by activation functions (i.e., $\mathbf{z}^{\ell}$). In contrast to other linear layers, these activated values undergo element-wise nonlinearity and are more likely to have independent meanings (Gurnee et al., 2023; Dar et al., 2023). Thus, without loss of generality, we zoom in and study monosemanticity based on one single layer of activated values, e.g., $\mathbf{z}^{\ell}$ in the $\ell$-th layer. Then, $f(\mathbf{W};\mathbf{x})$ can be divided into a frontal model $f_{1}(\cdot)$ and a followed model $f_{2}(\cdot)$ at $\ell$-th layer of output as follows:
(4) $\mathbf{z}=f_{1}(\mathbf{x})=\sigma^{\ell}(\mathbf{W}^{\ell}(\cdots\sigma^{2}(\mathbf{W}^{2}\sigma^{1}(\mathbf{W}^{1}\mathbf{x}))\cdots)),$
(5) $\mathbf{o}=f_{2}(\mathbf{z},\mathbf{x})=\text{Softmax}(\mathbf{W}^{L}\cdots\sigma^{\ell+2}(\mathbf{W}^{\ell+2}\sigma^{\ell+1}(\mathbf{W}^{\ell+1}\mathbf{z}))\cdots),$
where the frontal model $f_{1}(\cdot)$ takes the original data $\mathbf{x}$ as inputs and delivers output values in $\ell$-th layer $\mathbf{z}$ (with the superscript $\ell$ omitted for simplicity in notation.) As the rest components of the network, the following model $f_{2}(\cdot)$ takes $\mathbf{x}$ and $\mathbf{z}$ and outputs $\mathbf{o}$ (the entire model’s output), and $\mathbf{W}^{\ell}\in\mathbb{R}^{d_{\ell}\times d_{\ell+1}}$ denotes weights of linear transformation from the $\ell$-th layer to the $(\ell+1)$-th layer.
2.1. Activation and Monosemantic
Despite the progress of neural networks and emergence, there is no consensus definition for an “activated” neuron and a “monosemantic” neuron (Elhage et al., 2022). Here we try to summarize an intuitive introduction to describe them.
The Concept of Activation: If an input $\mathbf{x}^{[n]}$ triggers a neuron $z_{i}$ to output a value $f_{1}(\mathbf{x}^{[n]})_{i}$ that deviates “significantly” from the statistical mean (i.e., $\bar{z}_{i}$) , we say that neuron $z_{i}$ is activated by input $\mathbf{x}^{[n]}$. The challenge in defining activation lies in reaching a consensus on the meaning of “significantly”. Instead, we can provide a relative definition as follows. Generally, $i$-th neuron at $\ell$-th layer is considered to be more activated on $\mathbf{x}^{[2]}$ if:
(6) $\left\|\bar{z}_{i}-(f_{1}(\mathbf{x}^{[1]}))_{i}\right\|<\left\|\bar{z}_{i}-(f_{1}(\mathbf{x}^{[2]}))_{i}\right\|,$
[…]
Thus, to give a one-input-wise definition, one may rely on a threshold to define whether a neuron is activated, such as a hard threshold $\tau$. If the deviation of activation value from the mean $\|{\bar{z}_{i}-(f_{1}(\mathbf{x}^{[n]}))_{i}}\|$ exceeds $\tau$, it is generally considered that the neuron is activated. Otherwise, it is in an inactive state. The reason for setting $\tau$ is that neurons will generally have certain fluctuating outputs even for those unrelated different inputs. Unfortunately, setting a universal or adaptive $\tau$ remains to be explored. Therefore, it is also impractical to quantify activation by the threshold.
The Concept of Monosemanticity: However, further illustration for monosemanticity is closely related to a definition of activation that considers only one input at a time. We provide an abstract definition for it: $act(z_{i},\mathbf{x}^{[n]})$, which equals 1 when $z_{i}$ is activated by $\mathbf{x}^{[n]}$, and 0 otherwise. One can use a task-oriented definition as implementation.
To understand neural networks, an important research direction is to correlate neuron activations with human-interpretable features, such as Python and German for language processing; fur and grass for image processing, et al.. Existing works construct feature datasets $\{C_{1},\cdots,C_{m}\}$ for $m$ features, each containing a set of inputs. These feature datasets are specifically designed to partition the inputs $X=\{\mathbf{x}\}$, mathematically:
$\forall_{i\neq j}C_{i}\cap C_{j}=\emptyset;\bigcup_{i=1}^{m}C_{i}=X.$
A neuron $z_{i}$ is “monosemantic” if it is only activated on inputs that share a specific feature $C_{j}$, that is:
$\forall_{\mathbf{x}}act(z_{i},\mathbf{x})=1,\mathbf{x}\in C_{j}$
However, features are human-defined and vary a lot. In their study, Gurnee et al. (2023) consider about 100 features, while Trenton et al. (2023) study up to $10^{5}$ features to fully capture monosemanticity. Without unified feature datasets, it is also difficult to explicitly formalize the definition of “monosemantic”. More related works are discussed in Appx. B.
Thus, to detect monosemantic neurons, previous studies require manually labeled feature datasets and time-consuming offline computations after model training. To detect and inhibit monosemantic neurons during training, it is necessary to define a lightweight online metric $\phi(\cdot)$ that does not rely on feature datasets, which indicates the level of monosemanticity of a neuron.
2.2. Monosemanticity Inhibition
To achieve the desired output $z$, deep learning models use optimization strategy $\omega$ to update parameters ($\mathbf{W}$), such as minimizing the explicit loss function through gradient descent.
However, to inhibit monosemanticity, there is no related loss function to minimize. In this paper, we achieve this objective in two phases.
Recall that the model $f$ is parameterized by $\mathbf{W}$ and split into a frontal model $f_{1}$ and a followed model $f_{2}$ with respect to the studied layer of neurons $\mathbf{z}$.
Assume that we feed input $\mathbf{x}$ to $f$ and find $z_{i}\in\mathbf{z}$ is monosemantic. By using a well-designed optimization strategy $\omega$, parameters are updated to $\mathbf{W}^{*}$. By feeding the same input $\mathbf{x}$ into the neural network, the frontal model, the followed model, and the layer of neurons are updated to $f_{1}^{*}$, $f_{2}^{*}$, and $\mathbf{z}^{*}$, respectively. We hope that:
  • With updated $f_{1}$, neuron $z_{i}$ is less activated for input $\mathbf{x}$. Formally, given old $z_{i}\in\mathbf{z}=f_{1}(\mathbf{x})$ and updated $z_{i}^{*}\in\mathbf{z}^{*}=f_{1}^{*}(\mathbf{x})$, we expect:
    $\phi(z_{i}^{*})<\phi(z_{i}).$
@@ 2.2. Monosemanticity Inhibition
  • With updated $f_{2}$, the output is still robust when neuron $z_{i}$ is deactivated. Formally, replace $z_{i}$ with a weakly activated value $z_{i}^{\prime}$ (i.e., $\phi(z_{i}^{\prime})<\phi(z_{i})$) to obtain $\mathbf{z}^{\prime}$, we expect the loss $\mathcal{L}$ satisfies:
    $\mathcal{L}(f_{2}^{*}(\mathbf{z}^{\prime},\mathbf{x}))<\mathcal{L}(f_{2}(\mathbf{z}^{\prime},\mathbf{x})).$
@@
3. Methods
As discussed in Sec. 1, the existing neural network training paradigm is a passive process of decreasing monosemantic neurons. To proactively reduce monosemanticity, it is an intuitive solution to detect monosemantic neurons and inhibit them. Unfortunately, achieving these goals is challenging because the monosemanticity measurement does not exist and the simple inhibition method is counterproductive. Therefore, we first propose a new metric to measure the monosemanticity of artificial neurons with the guarantee of computation efficiency, then introduce a theoretically supported inhibition method to suppress monosemantic neurons to proactively reduce the ratios of monosemantic neurons in training artificial neural networks.
3.1. Metric for Monosemanticity
To proactively detect monosemantic neurons, it is important to design a metric that is general for different tasks and efficient to calculate. However, as discussed in Sec. 2, strictly defining “activated” is still under discussion in quantitative analysis (Elhage et al., 2022), which in turn leads to the difficulty of defining “monosemantic”. Although some pioneer works explore the measurement of monosemanticity, these metrics are inflexible for different tasks since they require a predefined and manually labeled feature dataset (Gurnee et al., 2023; Trenton et al., 2023). To address these limitations, we propose a robust metric $\phi(\cdot)$ that can accurately detect monosemantic neurons in this subsection. This metric fulfills two important criteria: (1) generality to make the metric do not rely on any specific dataset, (2) efficiency to enable fast online detection during training.
Intuitively, defining monosemantic neurons mainly requires starting from two principles: high deviation of activation value and low frequency of activation. First, a neuron is considered “activated” when its output value for the current input deviates from the mean value of outputs for all possible inputs. For a monosemantic neuron, its value distribution is more skewed and incurs large deviation (Elhage et al., 2022). Second, each monosemantic neuron only activates when its corresponding feature is inputted, which rarely happens considering the current datasets with steadily growing scales of samples and feature types. Following two principles, we formally define our metric Monosemantic Scale (MS for short) as follows.